# Dynamic link switching induces stable synchronized states in sparse networks

### v1

@@ Abstract
The flow of information in networked systems composed of multiple interacting elements strongly depends on the level of connectivity among these elements. Sparse connectivity often hinders the emergence of states where information is globally shared, such as fully synchronized states. In this context, dynamically switching the existing network links among the system elements can facilitate the onset of synchronization. However, determining the optimal rate at which links should be switched to ensure asymptotically stable synchronized behavior remains an open question. Here, we address this issue in a double-layer network of FitzHugh-Nagumo oscillators with sparse inter-layer connectivity. Specifically, we demonstrate that faster switching of the inter-layer links, which sweeps through a more significant number of oscillators, is more effective in establishing inter-layer synchronization than maintaining links on the identical oscillators for extended durations. We also evaluate the effectiveness of fast-switching links as the layer size increases. In a reduced system, we introduce smooth square-wave functions to emulate link switching, eventually introducing discontinuities into the system dynamics. This approach enables the assessment of the transverse stability of synchronized states induced by switching links.
@@ I Introduction
Many technological systems depend on efficient methods of information exchange between their various components in order to function properly. Examples include wireless sensor networks, where spatially distributed sensors communicate with central control systems [1]; smart grid systems, where data from distributed energy resources, substations, and smart meters converge at control centers [2]; and autonomous vehicle networks, where vehicles and infrastructure continuously exchange data to manage traffic, prevent collisions, and coordinate movement [3]. Interestingly, to minimize energy consumption at the component level and optimize bandwidth usage with the central control system, communication between components and the control center occurs intermittently, activated only when necessary to update the system. This intermittent exchange of information raises questions about the thresholds for the optimal duration that a connection must remain active to ensure consensus of information across the entire system.
In the theoretical study of dynamical systems, the effects of intermittent information exchange have been extensively explored in the context of synchronization in networks of coupled nonlinear systems. For instance, in a network of chaotic oscillators, it has been shown that coupling the oscillators for only a fraction of their state space can lead to synchronization, whereas permanent coupling within the full state space does not [4, 5]. Ref. [6] investigates the stability of synchronized states in networks where all links are switched on and off with specified time scales. Additionally, Ref. [7] demonstrates that chaotic oscillators periodically coupled using a sinusoidal function can achieve enhanced synchronization by appropriately adjusting the coupling frequency. In Ref. [8], two types of oscillators are intermittently coupled when the distances between specific state variables of certain oscillators fall below a predefined threshold. This approach revealed convergence to synchronization that depends on the chosen threshold. Furthermore, in a double-layer network of FitzHugh-Nagumo (FHN) oscillators, randomly switching the connections between oscillators in different layers induces an abrupt transition to inter-layer synchronization as the number of connections increases [9]. Finally, beyond the aforementioned studies, a wide range of other topological time dependencies have been investigated to enhance synchronization in both single-layer networks [10, 11, 12, 13, 14, 15] and multilayer networks [16, 17, 18, 19].
Despite advancements in the literature, little is known about the optimal time interval during which connections between certain elements should remain active before switching to different elements, ensuring that the entire system remains synchronized. Specifically, the following questions arise: given a limited number of connections, is it more advantageous for synchronization to maintain the connection between certain elements for a longer duration, or is it beneficial to shorten the connection duration and switch to different elements? Furthermore, are the synchronized states resulting from such switching asymptotically stable?
In this work, we investigate these questions within the framework of synchronization in a double-layer network of FitzHugh-Nagumo oscillators. Each isolated layer consists of two distinct populations of FHN oscillators diffusively coupled to their nearest neighbors via intra-layer connections. The individual dynamics of each layer is chaotic, with the oscillators within each layer remaining desynchronized. Additionally, the two identical layers are sparsely interconnected through inter-layer links, which diffusively couple pairs of oscillators from each layer (mirrored oscillators). The layers are initially desynchronized with respect to one another.
In this context, we periodically switch the existing inter-layer links at a fixed time interval, referred to as the switching time $T_{swt}$. At each switching event, the existing inter-layer links are randomly reassigned between corresponding oscillators in the two layers. We observe that shortening the switching time, thereby allowing inter-layer links to sweep through a larger number of oscillators across the layers, is more effective for achieving synchronization than maintaining the same pair of mirror oscillators connected for longer time intervals. This is demonstrated by analyzing the system’s time evolution under different switching times. Furthermore, the effectiveness of the switching procedure is evaluated for increasing layer sizes, highlighting its performance to establish synchronization across different switching times. Additionally, we investigate the time to synchrony for different switching times and population sizes, while keeping the ratio between the interlayer links and the population size constant, complying with the sparse network requirement. Also, the minimum switching time as a function of the population while keeping the ratio of the number of interlayer links to the population size constant for desynchrony is another approach to show, quantitatively, the synchronization dependency on the switching time and the population size. Finally, to examine the asymptotic stability of the synchronized states induced by the switching procedure, we analyze a reduced system consisting of two FHN oscillators per layer, interconnected by a single inter-layer link. To address the eventual discontinuities implied in the switching procedure, we model the switching of the inter-layer link using a smooth square-wave function. Using this approach, we apply the master stability function (MSF) formalism to determine the interval of switching times and coupling intensities for which the synchronized states are stable.
@@ II Model: A double-layer network of FitzHugh-Nagumo oscillators
To investigate the appropriate time intervals during which a network connection must remain active to ensure complete network synchronization, we consider a double-layer network with nodes composed of FitzHugh-Nagumo (FHN) oscillators. Each layer comprises $N$ FHN oscillators diffusively coupled to their nearest neighbors under periodic boundary conditions, forming ring topologies. The two layers are interconnected through a specified number of inter-layer links $N_{IL}$, which connect nodes positioned identically in each layer, referred to as mirror nodes (see the schematic picture in Fig. 1(a)). Multilayer networks have recently been employed to explore a variety of dynamical phenomena (e.g., [20, 21, 22, 23]); see also Ref. [24] for a comprehensive review. Similar multilayer arrangements have been extensively studied in both deterministic [25, 26, 27, 28, 29] and stochastic systems [30, 31, 32, 33]. Notably, the inter-layer coupling considered here is sparse, as the number of connected mirror nodes is small, i.e., $N_{IL}\ll N$. Additionally, in our double-layer network, the positions of the inter-layer links are time-dependent. These links remain active between a set of mirror nodes for a finite time interval before randomly switching to a different set. This interval is defined as the switching time, $T_{swt}$. To provide a more precise description of this system, we define its spatiotemporal dynamics using the following equations:
@@
where $\mathbf{x}_{i}=\left[u_{i1},v_{i1},\dots,u_{iN},v_{iN}\right]^{\top}$ represents the state vector of the $i$-th layer ($i=1,2$), which includes the activator $u_{ij}$ and inhibitor $v_{ij}$ of the $j$-th ($j=1,2,\dots,N$) node in the $i$-th layer. The dynamics of the FHN oscillators [34, 35] at the nodes of each layer are governed by:
$\mathbf{F}(\mathbf{x}_{i})=\begin{pmatrix}\frac{1}{\varepsilon}\left(u_{i1}-\frac{u_{i1}^{3}}{3}-v_{i1}\right)\\ u_{i1}+a_{i1}\\ \vdots\\ \frac{1}{\varepsilon}\left(u_{iN}-\frac{u_{iN}^{3}}{3}-v_{iN}\right)\\ u_{iN}+a_{iN}\end{pmatrix},$ (2)
where $\varepsilon=0.05$ is a control parameter that specifies the time scale separation between the fast activator $u_{ij}$ and the slow inhibitor $v_{ij}$. The excitability threshold for the parameter $a_{ij}$ determines whether the FHN oscillator operates in its excitable ($|a_{ij}|>1$) or oscillatory ($|a_{ij}|<1$) regime. All oscillators in this work operate in the oscillatory regime; however, each layer consists of two distinct populations of FHN oscillators, differentiated by their respective values of $a_{ij}$. These oscillators are alternately distributed throughout the ring topology: for odd-indexed oscillators ($j=1,3,5,\dots$), the value is $a_{ij}=0.87$, while for even-indexed oscillators ($j=2,4,6,\dots$), the value is $a_{ij}=0.97$ (see golden and blue nodes in Fig. 1(a)). This intra-layer heterogeneity reduces the likelihood of synchronization among the FHN oscillators within each layer. However, the two layers considered are identical, and each pair of mirror nodes hosts FHN oscillators of the same type, as illustrated in Fig. 1(a).
The parameters $\sigma_{1}$ and $\sigma_{2}$ in Eq. (1) regulate the coupling intensity among the FHN oscillators within layers $1$ and $2$, respectively. Additionally, the intra-layer nearest-neighbor coupling structure is defined by the zero row sum Laplacian matrix $\mathcal{L}$ as given by:
@@
where the phase angle $\phi\in[-\pi,+\pi)$ plays a crucial role in shaping the dynamic behavior within each isolated layer [26, 9]. Following the cited literature, we set the cross-coupling parameter to $\phi=\frac{\pi}{2}-0.1$, a value that allows desynchronized behavior within each layer.
The coupling between the layers (inter-layer coupling) is diffusive, with the intensity regulated by $\sigma_{12}$ and the structure defined by the inter-layer Laplacian $\mathcal{L}^{I}$. The inter-layer coupling matrix $\Gamma(t)$ governs the coupling exclusively between the activator variable $u$ of mirror oscillators. These matrices, $\mathcal{L}^{I}$ and $\Gamma(t)$, are defined as follows:
@@
The time dependence $(\gamma_{j}(t)$ in the components of $\Gamma(t)$ reflects the switching of inter-layer links at each time interval $T_{swt}$. If an inter-layer link exists for oscillator $j$, then $\gamma_{j}(t)=1$; otherwise, $\gamma_{j}(t)=0$. In contrast to [17], where a time-dependent rewiring is introduced implicitly, we apply a real time-dependent wiring, where we can define the rewireing frequency explicitly. Finally, the FHN oscillators in each layer are initialized with initial conditions (ICs) randomly selected from the intervals $u_{ij}(0)\in\left[-2.0,2.0\right]$ and $v_{ij}(0)\in\left[-2.0,2.0\right]$. In this way, both the FHN oscillators within the layers and the layers themselves are desynchronized at the initial time.
Following these definitions, we numerically integrate Eq. (1) to analyze the spatiotemporal dynamics of our double-layer network. As a first step, we consider each layer in isolation by setting the inter-layer coupling intensity to zero, i.e., $\sigma_{12}=0$. Meanwhile, the intra-layer coupling intensities are fixed at $\sigma_{1}=\sigma_{2}=0.1$. Each layer contains $N=200$ FHN oscillators. Figures 1(b) and 1(c) illustrate the spatiotemporal dynamics of the activator $u_{ij}$ for each FHN oscillator $j$ ($j=1,\dots,N$) in layers $i=1$ and $i=2$, respectively. These figures highlight the asynchronous dynamics both among the FHN oscillators within each layer and between the layers.
@@
Figure 1: (a) Schematic of a double-layer network illustrating the two distinct populations of FHN oscillators in its nodes, represented in blue and gold. The constants $\sigma_{1}$ and $\sigma_{2}$ denote the intra-layer coupling intensities, while $\sigma_{12}$ represents the inter-layer coupling intensity. (b) Spatiotemporal dynamics of the activator variable $u_{1j}$ for the FHN oscillators in layer $1$. (c) Spatiotemporal dynamics of the activator variable $u_{2j}$ for the FHN oscillators in layer $2$. Parameters are $\varepsilon=0.05$, $a_{ij}=0.87$ (even $j$), $a_{ij}=0.97$ (odd $j$), $\phi=\frac{\pi}{2}-0.1$, $\sigma_{12}=0$, $\sigma_{1}=\sigma_{2}=0.1$, and $N=200$.

### v2

@@ Abstract
The flow of information in networked systems composed of multiple interacting elements strongly depends on the level of connectivity among these elements. Sparse connectivity often hinders the emergence of states in which information is globally shared, such as fully synchronized states. In this context, dynamically switching existing network links among system elements can facilitate the onset of synchronization. Here, we address this problem in a double-layer network of FitzHugh–Nagumo oscillators with sparse inter-layer connectivity at fixed density. We show that dynamically switching the existing cross-layer links induces inter-layer synchronization, with a clear dependence on the switching time. In agreement with intuition, shorter switching times suppress large deviations between temporally connected oscillators and more effectively promote synchronization; crucially, this effect persists even when each isolated layer is chaotic. Chaos at the layer level is verified by a strictly positive largest Lyapunov exponent, confirming that synchrony is induced by switching rather than by periodic dynamics. For a minimal double-layer system, we emulate switching using smooth square waves and compute the master stability function (MSF), which is in agreement with direct numerical simulations and delineates the stability regions in parameter space.
@@ I Introduction
Many technological systems depend on efficient methods of information exchange between their various components in order to function properly. Examples include wireless sensor networks, where spatially distributed sensors communicate with central control systems [28]; smart grid systems, where data from distributed energy resources, substations, and smart meters converge at control centers [13]; and autonomous vehicle networks, where vehicles and infrastructure continuously exchange data to manage traffic, prevent collisions, and coordinate movement [42]. Interestingly, to minimize energy consumption at the component level and optimize bandwidth usage with the central control system, communication between components and the control center occurs intermittently, activated only when necessary to update the system. This intermittent exchange of information raises questions about the thresholds for the optimal duration that a connection must remain active to ensure consensus of information across the entire system.In the theoretical study of dynamical systems, the effects of intermittent information exchange have been extensively explored in the context of synchronization in networks of coupled nonlinear systems. For instance, in a network of chaotic oscillators, it has been shown that coupling the oscillators for only a fraction of their state space can lead to synchronization, whereas permanent coupling within the full state space does not [33, 40].Ref. [7] investigates the stability of synchronized states in networks where all links are switched on and off with specified time scales. Additionally, Ref. [15] demonstrates that chaotic oscillators periodically coupled using a sinusoidal function can achieve enhanced synchronization by appropriately adjusting the coupling frequency. Interestingly, periodic coupling may also be detrimental to synchronization [16, 2]. In Ref. [23], two types of oscillators are intermittently coupled when the distances between specific state variables of certain oscillators fall below a predefined threshold. This approach revealed convergence to synchronization that depends on the chosen threshold. Furthermore, in a double-layer network of FitzHugh-Nagumo (FHN) oscillators, randomly switching the connections between oscillators in different layers induces an abrupt transition to inter-layer synchronization as the number of connections increases [6]. Finally, beyond the aforementioned studies, a wide range of other topological time dependencies have been investigated to enhance synchronization in both single-layer networks [5, 46, 14, 47, 9, 12] and multilayer networks [29, 30, 4, 41].Despite advancements in the literature, little is known about the optimal time interval during which connections between certain elements should remain active before switching to different elements, ensuring that the entire system remains synchronized. Specifically, the following questions arise: given a limited number of connections, is it more advantageous for synchronization to maintain the connection between certain elements for a longer duration, or is it beneficial to shorten the connection duration and switch to different elements? Furthermore, are the synchronized states resulting from such switching asymptotically stable?In this work, we investigate these questions within the framework of synchronization in a double-layer network of FitzHugh-Nagumo oscillators. Each isolated layer consists of two distinct populations of FHN oscillators diffusively coupled to their nearest neighbors via intra-layer connections. The individual dynamics of each layer is chaotic, with the oscillators within each layer remaining desynchronized. Additionally, the two identical layers are sparsely interconnected through inter-layer links, which diffusively couple pairs of oscillators from each layer (mirrored oscillators). The layers are initially desynchronized with respect to one another.In this context, we periodically switch the existing inter-layer links at a fixed time interval, referred to as the switching time. At each switching event, the existing inter-layer links are randomly reassigned between corresponding oscillators in the two layers. We observe that shortening the switching time, thereby allowing inter-layer links to sweep through a larger number of oscillators across the layers, is more effective for achieving synchronization than maintaining the same pair of mirror oscillators connected for longer time intervals. This is demonstrated by analyzing the system’s time evolution under different switching times. Furthermore, the effectiveness of the switching procedure is evaluated for increasing layer sizes, highlighting its performance to establish synchronization across different switching times. Additionally, we investigate the time to synchrony for different switching times and population sizes, while keeping the ratio between the interlayer links and the population size constant, complying with the sparse network requirement. Also, the minimum switching time as a function of the population while keeping the ratio of the number of interlayer links to the population size constant for desynchrony is another approach to show, quantitatively, the synchronization dependency on the switching time and the population size. Finally, to examine the asymptotic stability of the synchronized states induced by the switching procedure, we analyze a minimal double-layer system consisting of two FHN oscillators per layer, interconnected by a single inter-layer link. To address the eventual discontinuities implied in the switching procedure, we model the switching of the inter-layer link using a smooth square-wave function. Using this approach, we apply the master stability function (MSF) formalism to determine the interval of switching times and coupling intensities for which the synchronized states are stable.
@@ II Model: A double-layer network of FitzHugh-Nagumo oscillators
To investigate the appropriate time intervals during which a network connection must remain active to ensure complete network synchronization, we consider a double-layer network with nodes composed of FitzHugh-Nagumo (FHN) oscillators. Each layer comprises $N$ FHN oscillators diffusively coupled to their nearest neighbors under periodic boundary conditions, forming ring topologies. The two layers are interconnected through a specified number of inter-layer links $N_{IL}$, which connect nodes positioned identically in each layer, referred to as mirror nodes (see the schematic picture in Fig. 1(a)). Similar double- and multilayer networks have recently been employed to explore a variety of dynamical phenomena (e.g., [19, 37, 36, 44]); see also Ref. [43] for a comprehensive review. Similar multilayer arrangements have been extensively studied in both deterministic [25, 31, 21, 32, 35] and stochastic systems [45, 17, 38, 18]. Here, we consider a fixed partial inter-layer link density $N_{IL}/N=0.25$. We emphasize that the total number of possible inter-layer links is N(N1)/2N(N-1)/2. Since $N_{IL}\ll N(N-1)/2$, the inter-layer connectivity considered here is indeed very sparse. Additionally, in our double-layer network, the positions of the inter-layer links are time-dependent. These links remain active between a set of mirror nodes for a finite time interval before randomly switching to a different set. This interval is defined as the switching time, $T_{\rm swt}$. To provide a more precise description of this system, we define its spatiotemporal dynamics using the following equations:
@@
where $\mathbf{L}_{i}=\left[u_{i1},v_{i1},\dots,u_{iN},v_{iN}\right]^{\top}$ represents the state vector of the $i$-th layer ($i=1,2$), which includes the activator $u_{ij}$ and inhibitor $v_{ij}$ of the $j$-th ($j=1,2,\dots,N$) node in the $i$-th layer. The dynamics of the FHN oscillators [8, 22] at the nodes of each layer are governed by:
$\mathbf{F}(\mathbf{L}_{i})=\begin{pmatrix}\frac{1}{\varepsilon}\left(u_{i1}-\frac{u_{i1}^{3}}{3}-v_{i1}\right)\\ u_{i1}+a_{i1}\\ \vdots\\ \frac{1}{\varepsilon}\left(u_{iN}-\frac{u_{iN}^{3}}{3}-v_{iN}\right)\\ u_{iN}+a_{iN}\end{pmatrix}$, (2)
where $\varepsilon=0.05$ is a control parameter that specifies the time scale separation between the fast activator $u_{ij}$ and the slow inhibitor $v_{ij}$. The excitability threshold for the parameter $a_{ij}$ determines whether the FHN oscillator operates in its excitable ($|a_{ij}|>1$) or oscillatory ($|a_{ij}|<1$) regime. All oscillators in this work operate in the oscillatory regime; however, each layer consists of two distinct populations of FHN oscillators, differentiated by their respective values of $a_{ij}$.These oscillators are alternately distributed throughout the ring topology: for odd-indexed oscillators ($j=1,3,5,\dots$), the value is $a_{ij}=0.87$, while for even-indexed oscillators ($j=2,4,6,\dots$), the value is $a_{ij}=0.97$ (see golden and blue nodes in Fig. 1(a)). This intra-layer heterogeneity reduces the likelihood of synchronization among the FHN oscillators within each layer. However, the two layers considered are identical, and each pair of mirror nodes hosts FHN oscillators of the same type, as illustrated in Fig. 1(a).The parameters $\sigma_{1}$ and $\sigma_{2}$ in Eq. (1) regulate the coupling intensity among the FHN oscillators within layers $1$ and $2$, respectively. Additionally, the intra-layer nearest-neighbor coupling structure is defined by the zero row sum Laplacian matrix $\mathcal{L}$ as given by:
@@
where the phase angle $\phi\in[-\pi,+\pi)$ plays a crucial role in shaping the dynamic behavior within each isolated layer [31, 6]. Following the cited literature, we set the cross-coupling parameter to $\phi=\frac{\pi}{2}-0.1$, a value that allows desynchronized behavior within each layer.The coupling between the layers (inter-layer coupling) is diffusive, with the intensity regulated by $\sigma_{12}$ and the structure defined by the inter-layer Laplacian $\mathcal{L}^{I}$. The inter-layer coupling matrix $\Gamma(t)$ governs the coupling exclusively between the activator variable $u$ of mirror oscillators. These matrices, $\mathcal{L}^{I}$ and $\Gamma(t)$, are defined as follows:
@@
The time dependence $\gamma_{j}(t)$ in the components of $\Gamma(t)$ reflects the switching of inter-layer links at each time interval, i.e., the switching time $T_{swt}$. If an inter-layer link exists for oscillator $j$, then $\gamma_{j}(t)=1$; otherwise, $\gamma_{j}(t)=0$. In contrast to [30], where a time-dependent rewiring is introduced implicitly, we apply a real time-dependent wiring, where we can define the rewireing frequency explicitly. Finally, the FHN oscillators in each layer are initialized with initial conditions (ICs) randomly selected from the intervals $u_{ij}(0)\in\left[-2.0,2.0\right]$ and $v_{ij}(0)\in\left[-2.0,2.0\right]$. In this way, both the FHN oscillators within the layers and the layers themselves are desynchronized at the initial time.Following these definitions, we numerically integrate Eq. (1) to analyze the spatiotemporal dynamics of our double-layer network. As a first step, we consider each layer in isolation by setting the inter-layer coupling intensity to zero, i.e., $\sigma_{12}=0$. Meanwhile, the intra-layer coupling intensities are fixed at $\sigma_{1}=\sigma_{2}=0.1$. Each layer contains $N=200$ FHN oscillators. Figures 1(b) and 1(c) illustrate the spatiotemporal dynamics of the activator $u_{ij}$ for each FHN oscillator $j$ ($j=1,\dots,N$) in layers $i=1$ and $i=2$, respectively. These figures highlight the asynchronous dynamics both among the FHN oscillators within each layer and between the layers.
@@
Figure 1: (a) Schematic of a double-layer network illustrating the two distinct populations of FHN oscillators in its nodes, represented in blue and gold. The constants $\sigma_{1}$ and $\sigma_{2}$ denote the intra-layer coupling intensities, while $\sigma_{12}$ represents the inter-layer coupling intensity. (b) Spatiotemporal dynamics of the activator variable $u_{1j}$ for the FHN oscillators in layer 11. (c) Spatiotemporal dynamics of the activator variable $u_{2j}$ for the FHN oscillators in layer 22. Parameters are $\varepsilon=0.05$, $a_{ij}=0.87$ (even $j$), $a_{ij}=0.97$ (odd $j$), $\phi=\frac{\pi}{2}-0.1$, $\sigma_{12}=0$, $\sigma_{1}=\sigma_{2}=0.1$, and $N=200$.
Due to the distinct set of initial conditions (ICs) assigned to each identical layer, their spatiotemporal dynamics, while uncoupled, as shown in Fig. 1, are desynchronized. When the layers are coupled through the diffusive inter-layer coupling defined in Eq. (1), synchronization may occur, typically depending on topological properties such as the coupling connectivity and strength or, as we demonstrate here, on the switching time. However, the intrinsic dynamical features of the layers are also decisive for the onset of synchronization. In particular, when the layer dynamics exhibit a limit cycle, synchronization is much easier to achieve because limit cycles possess strong transverse contraction, so that the inter-layer coupling only needs to align the phases. In contrast, when the layer dynamics are chaotic and characterized by positive Lyapunov exponents, synchronization requires the coupling to continuously suppress intrinsic exponential divergence, leading to narrower stability regions in the space of topological parameters (coupling intensity and connectivity). To confirm that, for the parameters considered here, the uncoupled layers are indeed chaotic, thereby posing greater challenges for the switching-time strategy, we compute in Fig. 2 the largest Lyapunov exponent of the identical layers using the Benettin method [3].
Figure 2: The largest Lyapunov exponent $\hat{\lambda}_{\max}(t)$ as a function of time for an isolated layer ($\sigma_{12}=0$), computed via the Benettin method (with periodic renormalization of tangent vectors), converging to $\lambda_{\max}\approx 0.04$. The simulation parameters are $\varepsilon=0.05$, $a_{\text{odd}}=0.87$, $a_{\text{even}}=0.97$, $\phi=\frac{\pi}{2}-0.1$, $\sigma_{1}=\sigma_{2}=0.1$, and $N=200$, with local coupling radius $r=1$. A positive $\lambda_{\max}$ confirms the presence of chaotic dynamics within each layer.